a bus is traveling at 50 mph then reduces its speed to 40 mph. the 220 mi. trip takes 5 hrs. how long does it travel at each speed? show equation.
Found 3 solutions by midwood_trail, Alan3354, drk:
Answer by midwood_trail(310)

A bus is traveling at 50 mph then reduces its speed to 40 mph. The 220 mile trip takes 5 hrs. How long does it travel at each speed? Show equation.
Let D = distance
Let r = rate = speed
Let t = time (in hours)
rt = D
Bus going at 50mph
50t = 220....This is the equation for the bus going at 50 mph.
To find the time for this first case, divide both sides by 50.
t = 220/50
t = 4.4 hours
If the travels at 50 mph, it will take the bus 4.4 hours to reach 220 miles.
===============================================================
For bus going at 40mph:
rt = D
40t = 220....This is the equation for the bus going at 40 mph.
Divide both sides by 40 to find the time it takes the bus to reach 220 miles at 40 miles per hour.
t = 220/40
t = 5.5 hours
Is this clear?




Answer by Alan3354(21605)

a bus is traveling at 50 mph then reduces its speed to 40 mph. the 220 mi. trip takes 5 hrs. how long does it travel at each speed? show equation.
---------------------
d = rt
t = d/r
x = distance at 50 mph
----------------
t = (x/50) + (220-x)/40 = 5 hours
Multiply thru by 200
4x + 5(220-x) = 1000
-x = -100
x = 100 miles at 50 mph --> 2 hours
--> 120 miles at 40 mph --> 3 hours

Answer by drk(1908)

here is the original; problem
"a bus is traveling at 50 mph then reduces its speed to 40 mph. the 220 mi. trip takes 5 hrs. how long does it travel at each speed? show equation."
This is an RTD problem. Here is a table based on the given information.
bus . . . . . . . . .rate . . . . . . . . .time . . . . . . . . . distance
. . . . . . . . . . . . 50 . . . . . . . . . . t . . . . . . . . .50t
. . . . . . . . . . . . 40. . . . . . . . . 5 - t . . . . . . . . 200 - 40t
. . . . . . . . . . . . . . . . . . . . . . . .5. . . . . . . . . 220
from column 3 we get
50t + 200 - 40t = 220
combine like terms and we get
10t + 200 = 220
subtract to get
10t = 20
t = 2
5 - t = 3
50 mph for 2 hours
40 mph for 3 hours